Sylva walked to Cathy's house at 2km/h, and then rode her bicycle back to her own house at 10km/h. The entire trip took 3 h. How long a time did Sylvia walk?
Answer by checkley71(8403)

X/2+X/10=3
(5X+X)/10=3
6X/10=3 NOW CROSS MULTIPLY.
6X=30
X=30/6
X=5 MILES FOR THE ONE WAY TRIP.
PROOF:
5/2+5/10=3
2.5+.5=3
3=3